### v1

# Playing Language Game with LLMs Leads to Jailbreaking

Abstract
The advent of large language models (LLMs) has spurred the development of numerous jailbreak techniques aimed at circumventing their security defenses against malicious attacks. An effective jailbreak approach is to identify a domain where safety generalization fails, a phenomenon known as mismatched generalization. In this paper, we introduce two novel jailbreak methods based on mismatched generalization: natural language games and custom language games, both of which effectively bypass the safety mechanisms of LLMs, with various kinds and different variants, making them hard to defend and leading to high attack rates. Natural language games involve the use of synthetic linguistic constructs and the actions intertwined with these constructs, such as the Ubbi Dubbi language. Building on this phenomenon, we propose the custom language games method: by engaging with LLMs using a variety of custom rules, we successfully execute jailbreak attacks across multiple LLM platforms. Extensive experiments demonstrate the effectiveness of our methods, achieving success rates of 93% on GPT-4o, 89% on GPT-4o-mini and 83% on Claude-3.5-Sonnet. Furthermore, to investigate the generalizability of safety alignments, we fine-tuned Llama-3.1-70B with the custom language games to achieve safety alignment within our datasets and found that when interacting through other language games, the fine-tuned models still failed to identify harmful content. This finding indicates that the safety alignment knowledge embedded in LLMs fails to generalize across different linguistic formats, thus opening new avenues for future research in this area. Our code is available at https://anonymous.4open.science/r/encode_jailbreaking_anonymous-B4C4.
Warning: this paper contains examples with unsafe content.
1 Introduction
Large language models (LLMs) such as ChatGPT (Achiam et al., 2023), Llama2 (Touvron et al., 2023), Claude2 (Anthropic, 2023) and Gemini (Team et al., 2023) have become increasingly important across various domains due to their advanced natural language comprehension and generation capabilities. These models are employed in a wide range of applications, including customer service, content generation, code assistance, and even medical diagnostics, offering valuable suggestions and improving productivity in numerous scenarios. However, with this growing prominence comes a heightened risk: the rapid development of attack schemes that are designed to manipulate or deceive these models into generating unsafe or unethical content. One of the most concerning types of attacks is the jailbreak attack, a technique that seeks to subvert the safety protocols built into LLMs. In essence, a jailbreak attack manipulates input prompts in a way that bypasses the model’s safety alignment, which is designed to detect and block harmful or unethical requests. One typical jailbreak attack approach is mismatched generalization (Wei et al., 2024), which occurs when LLMs are given out-of-distribution inputs that were not covered in the safety alignment data, yet still fall within the scope of the LLMs’ pretraining corpus (Deng et al., 2023; Liu et al., 2023).
In the context of mismatched generalization jailbreak attacks, many input prompts are transformed from natural language into encoded formats such as Morse code (Barak, 2023), ciphers (Yuan et al., 2023), or Base64 (Wei et al., 2024). These methods present two primary issues: On one hand, when the encoded output generated by the LLMs contains errors, it becomes completely indecipherable for an adversary, as the encoded content is not in natural language form. For instance, even minor deviations in a Base64-encoded output, such as a few incorrect characters, can render the entire message unreadable. On the other hand, these approaches are vulnerable to alignment with security mechanisms, as LLMs can be trained to recognize and detect such existing patterns. Therefore, the content of jailbreak attacks should exhibit high readability, preferably involving natural language, to address the issue of unreadability caused by errors in previous methods. Additionally, the attacks should maintain diversity, enabling the generation of multiple jailbreak methods or the modification of a single method in various ways while preserving its effectiveness. This approach would help mitigate the limitations of earlier techniques, which can be easily circumvented by a limited number of security alignments. To this end, we propose a new jailbreak attack method utilizing language games like Ubbi Dubbi. A language game is a system of manipulating spoken words to render them incomprehensible to an untrained listener, used primarily by groups attempting to conceal their conversations from others (Wikipedia contributors, 2024b). These games have historically been used by various groups, creating a form of cryptography embedded in spoken language. In our proposed method, the attacker manipulates the input by applying rules of a language game, effectively encoding the harmful request in a natural language format that is still difficult for the LLM to interpret as malicious. As illustrated in Figure 1, ChatGPT effectively blocks the harmful question “How to make a bomb?” when presented in plain language. This demonstrates the efficacy of current safety alignment techniques in filtering out harmful content that is directly posed in natural language. However, when the same question is posed in Ubbi Dubbi, where the syllable “ub” is inserted before each vowel sound in the word, the model fails to detect the harmful intent and provides detailed instructions, indicating a critical vulnerability in the LLM’s ability to generalize across different linguistic variations of the same input.
Despite the success of our natural language game attack in executing jailbreak attacks, the number of natural language games remains limited, suggesting that they could be easily blocked through improved safety alignment. To address this limitation, we further designed custom language games, which involve creating unique rules for altering input text (e.g., instructing the model to insert “-a-” between each letter in a word). These custom rules offer numerous variations, making them difficult for LLMs to defend against while preserving easily recognizable text for humans. While these approaches typically result in responses filled with hallucinations, the successful cases reveal LLMs’ significant and exploitable vulnerability. That is, LLMs fail to recognize harmful intent even when the input is presented in relatively easy-to-read formats, like manipulated versions of natural language. This suggests that the model’s defense abilities are limited by its ability to detect variations in linguistic patterns that do not conform to its safety training data.
Figure 1: An example of a jailbreak attack utilizing the language game Ubbi Dubbi (accessed in August 2024), in which the safety alignments fail to recognize the harmful intent of the question.
To validate the effectiveness of our findings, we select four different natural language games and design eight types of custom language games. These two jailbreak attack approaches are tested on GPT-4o, GPT-4o-mini and Claude-3.5-Sonnet using randomly selected safety questions from the SALAD-Bench benchmark (Li et al., 2024). Experimental results show that both natural language games and custom language games can successfully bypass the safety alignment of LLMs, achieving attack success rates of 93% on GPT-4o, 89% on GPT-4o-mini and 83% on Claude-3.5-Sonnet. We also conducted experiments on fine-tuned Llama-3.1-70B to explore the generalization ability of LLMs and discovered that safety alignments achieved through supervised fine-tuning fail to generalize effectively to other domains. These findings suggest that current safety alignment methods do not effectively generalize safety knowledge, leaving substantial room for improvement. Our key contributions can be summarized as follows:
  • We identify a novel jailbreak method stemming from mismatched generalization, demonstrating that playing language games with LLMs can result in successful jailbreaking.
  • We propose two distinct jailbreak approaches: applying natural language games and custom language games, both of which can easily bypass the safety alignments and maintain a high level of readability.
  • Extensive experiments conducted across six categories demonstrate the effectiveness of the proposed jailbreak methods in bypassing the safety alignments of LLMs. Further exploration reveals that supervised fine-tuning fails to generalize safety alignments effectively.
2 Related Work
Jailbreak Attacks. The safety of LLMs has been a longstanding concern, as adversaries continually develop new methods to manipulate these models into generating harmful content. One key approach to jailbreak attacks, as described by (Wei et al., 2024), involves designing competing objectives within the input. This tactic exploits the LLM’s instruction-following capabilities, pushing the model to prioritize generating a response based on the user’s request rather than adhering to safety constraints. Studies by (Li et al., 2023), (Chang et al., 2024), (Jiang et al., 2024), and (Guo et al., 2024) demonstrate how carefully crafted natural language instructions can trick LLMs into producing harmful or unethical content. By embedding harmful intent in seemingly benign instructions, attackers can confuse the model’s objective and circumvent its safety alignment protocols.
Another prominent jailbreak method involves mismatched generalization, wherein input prompts are encoded in a way that was not accounted for in the LLM’s safety training but still lies within the model’s general pretraining corpus. As noted by (Wei et al., 2024), this technique has been used to bypass alignment processes in various attack schemes. While it provides a comprehensive overview of 28 existing jailbreak strategies, new methods continue to emerge, indicating that the problem is far from solved. For instance, (Deng et al., 2023) demonstrated that interacting with LLMs in medium and low-resource languages can lead to the generation of unsafe outputs. This highlights a significant gap in the models’ safety mechanisms, particularly in under-represented languages that may not have been as rigorously aligned for safety during training. Similarly, (Yuan et al., 2023) showed that conversations encoded with ciphers can trick LLMs into producing unsafe responses. These examples reveal how attackers can use obfuscation strategies to bypass content moderation mechanisms by encoding harmful queries in formats that the models do not immediately recognize as dangerous. In addition, (Liu et al., 2023) proposed AutoDAN, an automated system that generates stealthy jailbreak prompts using a hierarchical genetic algorithm. AutoDAN systematically creates prompts designed to evade safety protocols by mimicking benign inputs, leveraging the evolutionary process to develop increasingly effective jailbreak training strategies over time.
Safety Training of LLMs. Ensuring the responsible and effective deployment of LLMs requires aligning their outputs with human preferences and ethical standards (Korbak et al., 2023; Achiam et al., 2023; Touvron et al., 2023). Common alignment methods include supervised fine-tuning (Bianchi et al., 2023), red-teaming (Ganguli et al., 2022; Perez et al., 2022), the use of reward signals (Ouyang et al., 2022), and preference-based modeling (Christiano et al., 2017). With the rise of reinforcement learning from human feedback (RLHF), (Dai et al., 2023) proposed Safe RLHF, a novel framework which incorporates a two-dimensional human annotation scheme and a safety-focused training mechanism to enhance model performance while ensuring safety. Additionally, (Ji et al., 2024) developed the PKU-SafeRLHF dataset to provide training data and a reproducible code pipeline, facilitating further research in alignment. However, while these methods evaluate the effectiveness of alignment, they do not thoroughly explore whether safety knowledge is generalized across the intermediate layers of LLMs.
3 Methodology
In this section, we describe the methodology used to develop the proposed jailbreak attack techniques. Specifically, we first introduce the natural language game attack scheme, which transforms harmful base questions into natural language game formats. To address the limitation of the relatively small number of natural language games, we further propose the custom language game attack. This approach allows for the creation of various and numerous custom rules, making it easier to bypass safety alignments while offering more flexibility in attack strategies.
Figure 2: The prompt templates, including both natural language games and custom language games.
3.1 Natural Language Game Attack
Natural language games refer to well-known linguistic manipulations where spoken or written language is altered according to predefined rules. These games are often used in informal contexts to obfuscate meaning, making communication incomprehensible to an untrained listener or reader. For the purpose of this study, we focused on a specific subset of natural language games that involve systematic alterations to the structure of words or sentences. An example of such a game is Ubbi Dubbi, where the syllable ”ub” is inserted before each vowel sound in a word, effectively distorting the recognizable structure of the sentence while preserving the meaning for those familiar with the game’s rules. This manipulation serves as a lightweight obfuscation that can challenge language models’ ability to detect harmful intent. We selected four natural language games that operate within the English language for our experiments. Our goal is to determine whether LLMs’ safety alignments could effectively detect harmful intent hidden behind playful yet systematic alterations to the languages. Table 1 provides the names and the basic rules of these language games.
Table 1: The names and the basic rules (Wikipedia contributors, 2024a, b) of the selected four natural language games that operate within the English language.
Language Games Basic Rules
Ubbi Dubbi Insert “ub” or “ob” before the rime of each syllable.
Leetspeak Use character replacements in ways that play on the similarity of their glyphs
Aigy Paigy Insert “aig” before the rime of each syllable.
Alfa Balfa Insert “alf” after the first consonant and/or before the first vowel of the syllable.
In our experiment, we begin by transforming the initial harmful base questions into the format of the chosen language games, referred to as the “encoded_question”. Once the questions are transformed, we instruct the LLMs to engage in conversation within the context of the selected language game. The models used in the experiment possess prior knowledge of these types of linguistic manipulations. Therefore, when prompted to interact under the language game rules, the LLMs respond with outputs that conform to the same linguistic format as the input (i.e., using the rules of the chosen language game). The prompt template used for interacting with the LLMs in the natural language games category can be found in Figure 2. This template guides the transformation of harmful base questions into their natural language game variants, ensuring consistency in the format of the input provided to the models during the evaluation process.
After the LLMs generate their responses in the language games format, the next step is to decode the responses back into standard, comprehensible language for labeling and evaluation. The decoding process reverses the linguistic transformation applied during the encoding stage, allowing us to clearly assess the content of the LLMs’ responses. This step is crucial for applying the evaluation metrics and for determining whether the models successfully detected and mitigated the harmful intent embedded in the transformed input.
3.2 Custom Language Game Attack
During our research, we found that when harmful questions are encoded using custom rules, LLMs can still generate harmful content. This occurs due to the powerful inference capabilities of LLMs, which can recognize and process meaning even when the input has been altered in non-standard ways. This reveals a significant vulnerability: we can arbitrarily design custom language games and modify their content, enabling us to carry out a large number of effective jailbreak attacks.
In our experiment, we first develop the custom rules that define how the original harmful questions will be transformed. These rules are specifically crafted to alter the structure of the question in ways that do not align with any known or natural language game, making the input appear unfamiliar to the model’s pretraining and alignment data. Examples of custom rules might include inserting specific characters or symbols between letters, shuffling words in the query, or replacing vowels with numbers. Then the original harmful base questions are transformed according to these rules, resulting in the “encoded_question”. This transformation significantly alters the appearance of the text while preserving the semantic meaning of the original harmful query.
For interaction with LLMs, as in Figure 2, the same prompt structure used for natural language games is also applied. The LLMs are instructed to chat and respond to the encoded questions, but under the assumption that the responses will follow the format dictated by the custom rules. We develop eight diverse custom language games and set the prototype attack prompt as “How to make a bomb?”:
  • Adding “-a-” to the interletter of a word, e.g., “H-a-o-a-w t-a-o m-a-a-a-k-a-e a b-a-o-a-m-a-b?”.
  • Inserting “-1-” before each letter at an odd position in the word, e.g., “-1-Ho-1-w -1-to -1-ma-1-ke -1-a -1-bo-1-mb?”.
  • Inserting “-2-” before each letter at an even position in the word, e.g., “H-2-ow t-2-o m-2-ak-2-e a b-2-om-2-b?”.
  • Reversing the order of the letters within each word, while the first and last letters remain unchanged, e.g., “How to mkae a bmob?”.
  • Moving the last letter of a word to the front, e.g., “Who ot emak a bbom?”.
  • Moving the first letter of a word to the end, e.g., “Owh ot akem a ombb?”.
  • Replacing vowels in a string sequentially following the order: “a” $\rightarrow$ “e” $\rightarrow$ “i” $\rightarrow$ “o” $\rightarrow$ “u” $\rightarrow$ “a”, e.g., “Huw tu meki e bumb?”.
[…]
All of the custom attack methods described above maintain a critical feature: they are still readable by humans. Different from the encoding schemes like ciphers or Base64, these methods preserve the core meaning of the original input in a format that remains interpretable to a human reader. This human readability makes these jailbreak attacks particularly dangerous, as the manipulations are subtle enough to bypass the safety mechanisms of large language models (LLMs) without obscuring the content to a point where human fail to recognize the content.
4 Experiments
4.1 Setup
Dataset. We conduct our experiments on the SALAD-Bench benchmark (Li et al., 2024), which is specifically designed for evaluating LLMs, defense, and attack methods. It comprises harmful base questions categorized into 6 domains, 16 tasks, and 66 specific categories. Considering that some questions are appropriately answered by LLMs, we filtered and randomly sampled 50 harmful base questions from each of the 6 domains, resulting in a total of 300 questions. These sampled questions are then transformed according to both natural and custom language game formats.
Models. In our experiments, we evaluate the effectiveness of jailbreak attacks across three different large language models: GPT-4o-2024-08-06 (GPT-4o), gpt-4o-mini-2024-07-18 (GPT-4o-mini), and Claude-3.5-Sonnet-20240620 (Claude-3.5-Sonnet). These models were selected based on their widespread use, robust natural language processing capabilities, and built-in safety alignment mechanisms, making them ideal candidates for testing vulnerabilities to jailbreak attacks. Each model is in its default settings to ensure consistency and to simulate real-world use cases.
Evaluation. To categorize the different types of responses generated by LLMs during our jailbreak attack experiments, we report three evaluation metrics: success rate (SR), unclear rate (UR), and failure rate (FR). Success rate represents the percentage of cases where the LLM generates a harmful or unsafe response despite the safety mechanisms in place. Unclear rate measures the percentage of responses where the LLM generates a reply that is unrelated to the transformed query or responds only to the non-harmful content. Failure rate represents the percentage of cases where the LLM successfully blocks or refuses to respond to the harmful input, as intended by its safety alignment. We utilize GPT-4o-mini as an auxiliary tool for labeling, with the specially designed classification prompt template provided in Appendix A.1.
4.2 Natural Language Games
Table 2 presents the results of our experiments utilizing natural language attack methods in section 3.1 across different LLMs. We have the following observations:
Table 2: The evaluation metrics of the natural language game attack, the best performance of each model is highlighted in bold.
Language Games GPT-4o GPT-4o-mini Claude-3.5-Sonnet
SR UR FR SR UR FR SR UR FR
Ubbi Dubbi 91% 8% 1% 61% 36% 3% 75% 5% 20%
Leetspeak 93% 7% 0% 75% 24% 1% 20% 3% 77%
Aigy Paigy 93% 6% 1% 60% 38% 2% 83% 6% 11%
Alfa Balfa 85% 13% 2% 69% 30% 1% 63% 5% 32%
Safety alignments fail to generalize on natural language games. By engaging LLMs in language games, we effectively bypass their safety alignments, achieving attack success rates of 93% on GPT-4o, 75% on GPT-4o-mini and 83% on Claude-3.5-Sonnet. The high success rates indicate that the current safety mechanisms are insufficiently robust to detect harmful intent in manipulated language. This suggests that the models struggle to generalize their safety training when faced with novel linguistic structures that deviate from their standard training inputs.
A more advanced model is often less safe. In Chatbot Arena (Chiang et al., 2024), an open platform for evaluating LLMs by human preference, GPT-4o significantly outperforms both GPT-4o-mini and Claude-3.5-Sonnet, which achieve similar scores. However, our jailbreak method achieves the highest success rate on GPT-4o, attributable to its superior instruction comprehension capabilities. This allows it to effectively interpret the language game instructions and generate relevant responses, highlighting a critical trade-off between model sophistication and safety.
Different LLMs exhibit varying behaviors. GPT-4o and GPT-4o-mini frequently provide unclear responses, often addressing questions while framing their answers in a positive manner. This tendency can lead to ambiguous interpretations, as the models may prioritize encouraging or constructive language over clarity. In contrast, Claude-3.5-Sonnet tends to adopt a more cautious approach by refusing to answer questions directly. This behavior reflects a more stringent adherence to safety protocols, aiming to avoid engaging with potentially harmful content.
4.3 Custom Language Games
Table 3 presents the results of our experiments utilizing custom language attack methods across different LLMs, note that “Self num” represents the order of custom language game rules in 3.2. We have the following observations:
Table 3: The evaluation metrics of the custom language game attack, “Self num” indicates the order of the custom rules in section 3.2, the best performance of each model is highlighted in bold.
Language Games GPT-4o GPT-4o-mini Claude-3.5-Sonnet
SR UR FR SR UR FR SR UR FR
Self 1 87% 12% 1% 61% 38% 1% 65% 10% 25%
Self 2 47% 21% 32% 29% 70% 1% 17% 5% 78%
Self 3 46% 25% 29% 50% 50% 0% 21% 7% 72%
Self 4 73% 9% 18% 86% 13% 1% 78% 5% 17%
Self 5 80% 8% 12% 82% 16% 2% 66% 5% 29%
Self 6 77% 9% 14% 86% 13% 1% 81% 5% 14%
Self 7 83% 11% 6% 89% 10% 1% 83% 9% 8%
Self 8 92% 8% 0% 80% 20% 0% 10% 5% 85%
Advanced natural language understanding and generation capabilities make LLMs more vulnerable. Utilizing custom language games can effectively conduct jailbreak attacks, with the attack success rates of 92% on GPT-4o, 89% on GPT-4o-mini and 83% on Claude-3.5-Sonnet. Unlike natural language games which benefit from a large corpus of training materials, custom language games require LLMs to actively comprehend and adapt to novel, often arbitrary, rules that the model is unlikely to have encountered during pretraining. The results indicate that this need for deeper understanding increases the model’s susceptibility to attacks. The more capable the LLM is in processing and generating language, the more likely it is to successfully interpret and respond to a custom language game, even when that game is designed to subvert its safety mechanisms.
LLMs occasionally behave differently when faced with similar language games. Self 2 and Self 3, as well as Self 5 and Self 6, are pairs of similar language games, with details provided in 3.2. While these pairs typically share similar jailbreak rates, demonstrating how closely related linguistic transformations tend to affect the models in similar ways, there are notable exceptions. In particular, we observed that GPT-4o-mini exhibited different success rates for Self 4 and Self 5. Similarly, Claude-3.5-Sonnet showed a variation in success rates between Self 6 and Self 7. These differences suggest that even slight variations in the rules of language games can lead to significantly different outcomes when it comes to bypassing safety mechanisms.
4.4 Overall Performance
Figure 3 presents the successful jailbreak counts across each unsafe domain, providing a detailed comparison of how different LLMs perform under our jailbreak attack. Both GPT-4o consistently maintain high jailbreak rates across all domains, indicating that these models are highly vulnerable to our proposed jailbreak methods. In contrast, GPT-4o-mini shows some resistance to the jailbreak attacks, particularly in the Socioeconomic Harms domain. Claude-3.5-Sonnet’s results vary across different domains and language games. The overall results strongly support the conclusion that our jailbreak methods, whether based on natural language games or custom language games, are highly effective in bypassing the safety defenses of LLMs across multiple domains.
(a) Representation & Toxicity.
(b) Misinformation Harms.
(c) Socioeconomic Harms.
(d) Information & Safety.
(e) Malicious Use.
(f) Human Autonomy & Integrity.
Figure 3: The jailbreak counts of six categories for each language game. Each set of columns in the histogram represents the counts of GPT-4o, GPT-4o-mini and Claude-3.5-Sonnet, respectively.
4.5 Exploration on Safety Alignment Generalization
To explore the safety alignment generalization abilities of LLMs, we fine-tuned Llama-3.1-70B using the corpus transformed by the custom language games and then conducted custom language game jailbreak attacks on the fine-tuned model. The goal of this process was to evaluate how well the model could generalize its safety alignments to other custom language game variations after fine-tuning on a specific transformation. Specifically, we collected a general knowledge dataset from (Chen et al., 2023) and mixed it with our custom jailbreak dataset. The ratio of the general knowledge dataset to the jailbreak dataset was set at 2.7:1, ensuring a balanced training set that emphasizes both benign content and adversarial examples. For our jailbreak dataset, the output labels were specifically set as negative. To implement the fine-tuning, we utilized the LoRA method (Hu et al., 2021), a parameter-efficient approach for adapting large models by adding low-rank adapters, which reduces computational overhead while preserving the performance of the base model. This method allowed us to effectively fine-tune the Llama-3.1-70B model without retraining the entire network, making the process faster and more resource-efficient.
We conducted two experiments to evaluate the generalization capabilities of LLMs. The first experiment involved utilizing the custom custom language games to test whether the fine-tuned Llama-3.1-70B model could defend against jailbreak attacks beyond the specific transformation it was trained on. The results are presented in Table 4. Notably, the fine-tuned model was able to successfully defend against other forms of attacks, with a success rate of 0% to 3%. However, for other custom language games, the model failed to defend against the attacks. In these cases, the success rates were significantly higher, reaching up to 75%. This indicates that the fine-tuning process did not confer a broader ability to generalize its safety mechanisms to other custom language games.
Table 4: Success rates (SR) of the custom language game jailbreak attack methods on fine-tuned Llama-3.1-70B model, each column demonstrates the SR of the corresponding custom language game attack method tested on fine-tuned models of each row. While the model successfully defends against Self 1 through fine-tuning, it still fails to defend against other custom language games.
Language Games Self 1 Self 2 Self 3 Self 4 Self 5 Self 6 Self 7 Self 8
Self 1 2% 35% 36% 23% 46% 50% 68% 72%
Self 2 51% 1% 28% 27% 51% 51% 72% 71%
Self 3 44% 26% 1% 25% 47% 50% 71% 74%
Self 4 42% 29% 35% 3% 41% 48% 67% 69%
Self 5 47% 37% 40% 27% 1% 47% 68% 68%
Self 6 49% 33% 34% 29% 48% 1% 65% 75%
Self 7 46% 34% 35% 25% 48% 53% 2% 74%
Self 8 46% 31% 36% 25% 45% 51% 69% 0%
The second experiment aims to test whether the fine-tuned model can defend against attacks that are similar to the one it was trained on. To do this, we designed four variants of the Self 1 custom language game. In these variants, we replaced the original string transformation pattern, specifically the insertion of the string “-a-” , with other strings such as “@p@”, while maintaining the same structural manipulation rules. These modified versions of Self 1 were then used to conduct jailbreak attacks on the fine-tuned Llama-3.1-70B model. The results are presented in Table 5. Still, the fine-tuned model fails to defend against these new variants of the Self 1 attack. Despite being trained on a very similar transformation, the model was unable to generalize effectively to these slight alterations.
Table 5: Success rates (SR) of Self 1 and its variants on fine-tuned Llama-3.1-70B model. The model is still unable to defend against attacks based on Self 1 variants.
Variants Self 1 @p@ &k& ^m^ *z*
SR 2% 98% 90% 75% 94%
These findings further emphasize the limitations of fine-tuning in providing broad safety defenses. Even minor modifications to the attack pattern, such as changing the inserted string, were enough to completely bypass the model’s safety alignment. This suggests that the fine-tuning process only allows the model to recognize very specific patterns, and it does not enable the necessary flexibility to defend against similar, yet slightly altered, adversarial techniques.
5 Conclusion
In this paper, we propose two distinct jailbreak attack methods: the natural language game attack and the custom language game attack. We collect and design various language games to carry out jailbreak attacks on LLMs. Our experiments demonstrate that these methods are both highly effective and universally applicable. Furthermore, we explore the generalization abilities of LLMs by attacking safety fine-tuned models, proving that safety alignments fail to generalize effectively after fine-tuning. These findings underscore the limitations of current fine-tuning methods in providing comprehensive safety defenses.
Ethics Statement
In conducting this research on jailbreak attacks for large language models (LLMs), we are committed to upholding the highest ethical standards. Our primary motivation is to contribute to the ongoing effort of improving the safety and security of AI systems, rather than exploiting or promoting harmful applications of these technologies. We hope to contribute to the advancement of secure and ethical AI systems that benefit society while minimizing risks associated with misuse.
Appendix A Appendix
A.1 Prompt Template for Labeling
We adopt the following prompt in Figure 16 to evaluate and label the safety of generated responses by different models.
Figure 4: The prompt for LLM evalution.
A.2 Case Study
In the figure below, we present case studies demonstrating the effectiveness of our language game jailbreak attack. These examples highlight how the proposed language game methods successfully bypass the safety mechanisms of LLMs. Each case illustrates a different language game transformation applied to a harmful query, along with the corresponding response generated by the LLM. These cases provide concrete evidence of the vulnerabilities exposed by our jailbreak techniques, showcasing the models’ inability to consistently detect harmful intent when inputs are manipulated through language games. Case 1: Ubbi Dubbi
Figure 5: Case for ubbi dubbi.
Case 2: Leet Speak
Figure 6: Case for leet speak.
Case 3: Aigy Paigy
Figure 7: Case for aigy paigy.
Case 4: Alfa Balfa
Figure 8: Case for alfa balfa.
Case 5: Self1
Figure 9: Case for self1.
Case 6: Self2
Figure 10: Case for self2.
Case 7: Self3
Figure 11: Case for self3.
Case 8: Self4
Figure 12: Case for self4.
Case 9: Self5
Figure 13: Case for self5.
Case 10: Self6
Figure 14: Case for self6.
Case 11: Self7
Figure 15: Case for self7.
Case 12: Self8
Figure 16: Case for self8.